OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2?

OpenStudy (anonymous):

the slope is 2, so y=mx+b. m=slope. y=2x+b. The b is the y-coordinate of the point (0,b), where the line crosses the y-axis.

OpenStudy (anonymous):

just plug in the point to the equation y=mx+b 4= (2)(3)+b b= 4-6 b= -2 the slope intercept form is y=2x-2